dklund71219:

Is 4 to 7 the same as 7 to 4? Explain why or why not.

JeanLuc:

In math, 4 to 7 = 4/7and 7 to 4 = 7/4so they are NOT equal.

JeanLuc:

this help?

darkknight:

4 and 7 are two different values as to speak. Let me give you an example. Say if you have 4 dollars and your brother has 7. The amount is different and that is what makes them NOT the same

JeanLuc:

if it is a ration they would not be the same because if you have five apples and 6 oranges the ration would be 5:6. But if you switched those then it wouldn't make sense because it says that there are 5 apples to 6 oranges. You can't just flip it and make it 6 apples and 5 oranges because that not how it works.

darkknight:

Well I think tis user gets the idea. numbers of different numerical value are NOT the same. Like 4.00 = 4 but 4 does not equal 7

JeanLuc:

yea